OpenStudy (anonymous):

for which real numbers a, b, c will the function f(x)=(ax+b)/(cx+d) satisfy f(f(x))=x for all x

OpenStudy (anonymous):

Expand the f(f(x)) and try to find something

OpenStudy (anonymous):

What I got is a=cx and b=dx

OpenStudy (anonymous):

Keep the quotient form, you will have\[\frac{ a \frac{ ax+b }{ cx+d } +b }{ c \frac{ ax+b }{ cx+d }+ d } =x\]

OpenStudy (anonymous):

i know,then i get\[-x^{2}(ac+bc)-x(db-a^{2})+b(a+b)=0\], but what i can do next?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[(a-cx)\frac{ ax+b }{ cx+d }+(b-dx)=0\]

OpenStudy (anonymous):

then a=cx and b=dx is a solution to this equation

OpenStudy (asnaseer):

I believe you should multiply both sides by (cx + d) and then simplify first.

OpenStudy (asnaseer):

What I end up with is:\[c(a+d)x^2+(d^2-a^2)x-b(a+d)=0\]

OpenStudy (asnaseer):

Since this has to hold for ANY x, then every term in this equation must be zero.

OpenStudy (asnaseer):

The simplest solution is therefore:\[a=-d\]

OpenStudy (anonymous):

why every term in your equation must be zero?

OpenStudy (asnaseer):

because it has to hold for ANY x

OpenStudy (asnaseer):

i.e. for f(f(x)) = x for all x, we must have:\[c(a+d)x^2+(d^2-a^2)x-b(a+d)=0\]

OpenStudy (asnaseer):

for all x

OpenStudy (anonymous):

got it

OpenStudy (asnaseer):

this therefore leads to:\[f(x)=\frac{ax+b}{cx-a}\]

OpenStudy (asnaseer):

for any a, b and c will give:\[f(f(x))=x\]

OpenStudy (anonymous):

yeah, i got it, thx a lot

OpenStudy (asnaseer):

yw :)